95. What is the sustainable growth rate assuming the following ratios are constant? 1.46 8 percent 1.20 32 percent Total asset turnover Profit margin Equity multiplier Payout ratio A. 10.30 percent B. 10.53 percent C. 10.67 percent D. 10.89 percent E. 11.01 percent
Answer #1

ROE = Total asset turnover * Profit margin * Equity Multiplier =1.46* 8%*1.20 = 14.016%

Growth = ROE*(1-Payout Ratio)/(-(1-Payout ratio)*ROE) = 14.016%*(1-32%)/(1-14.016%*(1-32%)) = 10.53%

Option b is correct

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well